# Bridging the Gap: Protocol Towards Fair and Consistent Affect Analysis

### v1

@@ I Introduction
Automatic affect analysis has a long history of studies in the intersection of physiology, psychology, and machine/deep learning [25, 54, 20, 33, 31, 3, 5, 11]. The analysis involves automatic: a) recognition of categorical affect, represented via the so-called six basic expressions (e.g., anger, disgust, happiness), plus the neutral state [9]; b) detection of activation of facial action units (AUs) (i.e., specific movements of facial muscles) [9]; c) estimation of dimensional affect, represented via the emotion descriptors of valence (characterises an emotional state on a scale from positive to negative) and arousal (characterises an emotional state on a scale from active to passive) [56] [21, 36, 30, 57, 24, 23, 41].
Affect analysis methods are being developed using existing databases with the goal of maximizing their performance on the test sets and surpassing the performance of other methods [18, 35, 26, 17, 19, 4]. In most of the cases, this is the only criterion and there is no checking whether these comparisons are fair and/or whether the methods are unbiased and perform equally well across subjects of various demographic attributes such as race, gender and age (especially considering that the utilized databases do not contain an even distribution of subjects among these groups). The methods, unless explicitly modified, are severely impacted by such bias since they are given more opportunities (training samples) for optimizing their objectives towards the majority demographic group in the database. This leads to lower performance for the minority groups, i.e., subjects represented with less number of samples, which is not desired.
In the past three decades, many affective databases have been constructed with a clear tendency from small-scale to large-scale and from lab-controlled to real-world unconstrained conditions (termed “in-the-wild”) [67, 38, 27, 40, 39, 29, 15, 32, 17, 33, 16, 34, 28, 37, 22]. Two of the most typically utilized in-the-wild databases for Facial Expression Recognition (FER) are RAF-DB [43] and AffectNet [51]. AffectNet is further annotated in terms of valence and arousal (VA). The most widely used databases for AU recognition are DISFA [49], GFT [12] and the in-the-wild EmotioNet [6] and RAF-AU [65]. All these databases have some problems that we present next.
@@ II-A Annotation
We annotated all affective databases (AffectNet, RAF-DB, DISFA, EmotioNet, GFT and RAF-AU) in terms of demographic attributes of age, gender, and race. Let us note that although, in practice, race and ethnicity terms are used interchangeably, they are not the same. Race and ethnicity are different categorizations of humans. Race is defined based on physical traits and ethnicity is based on cultural similarities [59]. Therefore we decided to annotate the databases in terms of race (facial appearance is easier to judge and annotate). We adopted a commonly accepted race classification from the U.S. Census Bureau and thus we defined the following 5 race groups (in alphabetical order): Asian, Black (or African American), Indian (or Alaska Native), Native Hawaiian (or Other Pacific Islander), and White (or Caucasian). During the annotation process, we did not find any cases of Hawaiian or Other Pacific Islanders.
@@ II-D Performance Measures
In the following, we present the performance measures that we selected for validating the performance and fairness of the models on each task (Expression Recognition, AU Detection and VA Estimation).

### v2

@@ I Introduction
Automatic affect analysis has a long history of studies in the intersection of physiology, psychology, and machine/deep learning. The analysis involves automatic: a) recognition of categorical affect, represented via the so-called six basic expressions (e.g., anger, disgust, happiness), plus the neutral state [6]; b) detection of activation of facial action units (AUs) (i.e., specific movements of facial muscles) [6]; c) estimation of dimensional affect, represented via the emotion descriptors of valence (characterises an emotional state on a scale from positive to negative) and arousal (characterises an emotional state on a scale from active to passive) [24].
Affect analysis methods are being developed using existing databases with the goal of maximizing their performance on the test sets and surpassing the performance of other methods. In most of the cases, this is the only criterion and there is no checking whether these comparisons are fair and/or whether the methods are unbiased and perform equally well across subjects of various demographic attributes such as race, gender and age (especially considering that the utilized databases do not contain an even distribution of subjects among these groups). The methods, unless explicitly modified, are severely impacted by such bias since they are given more opportunities (training samples) for optimizing their objectives towards the majority demographic group in the database. This leads to lower performance for the minority groups, i.e., subjects represented with less number of samples, which is not desired.
In the past three decades, many affective databases have been constructed with a clear tendency from small-scale to large-scale and from lab-controlled to real-world unconstrained conditions (termed “in-the-wild”). Two of the most typically utilized in-the-wild databases for Facial Expression Recognition (FER) are RAF-DB [12] and AffectNet [20]. AffectNet is further annotated in terms of valence and arousal (VA). The most widely used databases for AU recognition are DISFA [18], GFT [8] and the in-the-wild EmotioNet [3] and RAF-AU [32]. All these databases have some problems that we present next.
@@ II-A Annotation
We annotated all affective databases (AffectNet, RAF-DB, DISFA, EmotioNet, GFT and RAF-AU) for demographic attributes of age, gender, and race. Let us note that although, in practice, race and ethnicity terms are used interchangeably, they are not the same. Race and ethnicity are different categorizations of humans. Race is defined based on physical traits and ethnicity is based on cultural similarities [26]. Thus we decided to annotate the databases in terms of race (facial appearance is easier to judge and annotate). We adopted a commonly accepted race classification from the U.S. Census Bureau; we defined the following 5 race groups (in alphabetical order): Asian, Black (or African American), Indian (or Alaska Native), Native Hawaiian (or Other Pacific Islander), and White (or Caucasian). During the annotation process, we did not find any cases of Native Hawaiian or Other Pacific Islander.
@@ II-D Performance Measures
In the following, we present the performance measures that we selected for validating the performance and fairness of the models on each task.